OpenStudy (anonymous):

Using the following system of inequalities find the maximum value of f(x,y) = 3x + 3y x≥0 y≥0 3x+2y≤18 6x+7y≤42

OpenStudy (anonymous):

@Stewpeas_0831 help me please

OpenStudy (anonymous):

@UsukiDoll help?

OpenStudy (anonymous):

Use an online calculator called mathway. Heres the link ;) https://mathway.com/ Sorry I couldn't teach you ;(

OpenStudy (anonymous):

can you do it . it wont allow me on the site

OpenStudy (anonymous):

3x + 2y = x = 6 + 0.6666666667y

OpenStudy (loser66):

The first equation: \(3x+2y\leq 18\), Can you graph it out??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Jessica did you get the answer??? I'm seeking it too. No matter how I work this I can't find the solution.